OpenStudy (anonymous):

Can some one tell me the formulas to find the following for quarter, nickel and dime. Formulas needed are: 1. area 2. surface area 3.volume (if any) With given diameters. Medals! @eSpeX

OpenStudy (anonymous):

@amistre64

OpenStudy (espex):

If you have the diameter, then you know the radius, all of these coins are cylinders or circles depending on the depth of information you are asked for.

OpenStudy (anonymous):

they are cylinders? I was thinking maybe, but was not sure.

OpenStudy (anonymous):

I am just to find the info I type above for each. had trouble finding it on line.

OpenStudy (anonymous):

So I can use the diameter measure to find the info above?

OpenStudy (espex):

Well the relative formulas are: \[Cylinder~Volume = \pi r^2h\]\[Circle~Area=\pi r^2\]\[Cylinder~Surface~Area=2\pi r^2+2\pi r h\]\[Area~of~Sides=2\pi rh\]

OpenStudy (espex):

Well you know that the diameter is equal to 2r

OpenStudy (anonymous):

ok

OpenStudy (espex):

So you can then get the area of one of the faces, times that by 2 and you have the area of both faces. If you assume no height, then you have no volume, and you only need to calculate one of the faces as it would be a 2D shape.

OpenStudy (anonymous):

So I find the area first to find the volume next right?

OpenStudy (espex):

You will need to get a height to find volume.

OpenStudy (anonymous):

ok I will get on it. Thank you again!

OpenStudy (espex):

You're welcome.